# On the realization of infinite (continuous)

### v2

@@ 1 Introduction
One of the relatively recent trends in the higher spin field theory [^1] Current state of the higher spin field problem is considered, e.g., in the reviews [1, 2, 3]. is the study of the Lagrangian formulation of infinite (continuous) field models in $\rm{AdS}$ space. Various approaches to such a formulation are discussed in refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. [^2] The review of the infinite (continuous) spin field models is given in ref. [14], for more recent references see e.g., [12]. In particular, a full classification of possible unitary and non-unitary continuous spin field models was given in the paper by R. Metsaev [9] within the light-cone formalism.
@@
In a recent paper [12], a completely consistent covariant Lagrange formulation for bosonic infinite spin field in $\rm{AdS}_{4}$ space was constructed in the framework of the BRST approach to higher spin fields. The formulation is realized in the Fock state of the vectors $|{\Psi}\rangle$ which are annihilated by a set of operator constraints satisfying the closed algebra (see relations (2.4) bellow). The Lagrangian includes the real continuous parameter $\boldsymbol{\mu}$ and in the flat limit this Lagrangian is reduced to the known infinite spin field Lagrangian in Minkowski space in which $\boldsymbol{\mu}$ defines the corresponding Poincaré representation. This allows us to treat the vector $|{\Psi}\rangle$ as an infinite spin field in $\rm{AdS}_{4}$ space. Then natural questions arise as to what representation of the symmetry group of $\rm{AdS}$ space such a formulation corresponds and how it relates to Metsaev’s classification. In this paper we answer these questions. First, we show that the field $|{\Psi}\rangle$ in the paper [12] corresponds to the most degenerate continuous unitary representation of the Lie algebra $\mathfrak{so}(2,3)$. [^3] The notion of the most generate representation is based on the Gelfand theorem [15] which states that the number of independent invariant operators in the enveloping algebra acting in the Hilbert space of functions on the domain $X$ can be independent of the rank of the group. The most degenerate representation, if it exists in a given space, means that the number of these operators is one. The construction of the most degenerate continuous representations of the group $\mathrm{SO}(p,q)$ is discussed in the papers [16, 17, 18], see also [19]. Second, we show that the representations of the $\mathrm{SO}(2,3)$ group realized on the field $|{\Psi}\rangle$ correspond to the $\mathrm{SO}(2,3)$ representations in Metsaev’s classification. In this case, the eigenvalues of the $\mathfrak{so}(2,3)$ Casimir operators are determined by the dimensionless quantity $\boldsymbol{\mu}R$ where $R$ is the radius of $\rm{AdS}_{4}$ space.
@@
The paper is organized as follows. In Section 2, we briefly review the basic results of the paper [12]. Here the infinite spin field in $\rm{AdS}_{4}$ space is defined, the operator constraints $L_{0},L,L^{+}$ are introduced, their algebra is written down and the flat limit is discussed. In Section 3, the geometric structure of $\rm{AdS}_{4}$ space is described in a form convenient for further consideration. In Section 4, we show that the operator constrains are the invariants of the $\mathrm{SO}(2,3)$ group and a more invariant operator $(\Pi)^{2}$ is introduced. In Section 5, we construct the Casimir operators $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ of the $\mathfrak{so}(2,3)$ algebra acting in the Fock space of fields $|{\Psi}\rangle$ under the conditions (2.4). In doing so, we have obtained two important statements. First, we proved that $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ are completely expressed in terms of the operators of the constraints (2.4) under consideration. Second, we show that only one of two these Casimir operators is independent. This exactly corresponds to the definition of the most degenerate representation for the $\mathrm{SO}(2,3)$ group. The summary is a collection of results.
@@ 2 Infinite spin field in $\rm{AdS_{4}}$ space
We begin with a brief summary of the basic results of the paper [12] on the Lagrangian formulation of the infinite spin field in $\rm{AdS_{4}}$ space.
The Lagrangian formulation under consideration was proposed within the BRST approach to higher spin field theory in terms of the Fock space with the vectors
$|{\Psi}\rangle=\sum\limits_{s_{1},s_{2}=0}^{\infty}\frac{1}{(s_{1}!s_{2}!)^{1/2}}\,\psi_{\alpha_{1}\ldots\alpha_{s_{1}}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s_{2}}}(x)c^{\alpha_{1}}\ldots c^{\alpha_{s_{1}}}\bar{c}_{\dot{\alpha}_{1}}\ldots\bar{c}_{\dot{\alpha}_{s_{2}}}|0\rangle\,,$ (2.1)
where $\psi_{\alpha_{1}\ldots\alpha_{s_{1}}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s_{2}}}(x)$ are the fields in $\rm{AdS_{4}}$ space with the coordinates $x^{\mu}$, $\mu=0,1,2,3$. Here $a_{\alpha}$, $\bar{a}^{\dot{\alpha}}$ and $c^{\alpha}=(\bar{a}^{\dot{\alpha}})^{\dagger}$, $\bar{c}_{\dot{\alpha}}=(a_{\alpha})^{\dagger}$ are the annihilation and creation operators, respectively. These operators satisfy the following commutation relations: [^4] The two-component Weyl spinor indices are raised and lowered by $\epsilon_{\alpha\beta}=-\epsilon^{\alpha\beta}$, $\epsilon_{\dot{\alpha}\dot{\beta}}=-\epsilon^{\dot{\alpha}\dot{\beta}}$: $\varphi^{\alpha}=\epsilon^{\alpha\beta}\varphi_{\beta}$, $\bar{\varphi}^{\dot{\alpha}}=\epsilon^{\dot{\alpha}\dot{\beta}}\bar{\varphi}_{\dot{\beta}}$. The relativistic Pauli matrices $\sigma^{m}_{\alpha\dot{\beta}}$ and $\tilde{\sigma}_{m}^{\dot{\alpha}\beta}$ satisfy the basic relation $\sigma_{m}\tilde{\sigma}_{n}+\sigma_{n}\tilde{\sigma}_{m}={}-2\eta_{mn}$ where $\eta_{mn}=\mathrm{diag}(-1,+1,+1,+1)$. We use the matrices $\sigma_{mn}=-\left(\sigma_{m}\tilde{\sigma}_{n}-\sigma_{n}\tilde{\sigma}_{m}\right)/4$, $\tilde{\sigma}_{mn}=-\left(\tilde{\sigma}_{m}\sigma_{n}-\tilde{\sigma}_{n}\sigma_{m}\right)/4$. Here and further the Latin indices take the values 0,1,2,3.
$[a_{\alpha},c^{\beta}]=\delta_{\alpha}^{\beta}\,,\qquad[\bar{a}^{\dot{\alpha}},\bar{c}_{\dot{\beta}}]=\delta_{\dot{\beta}}^{\dot{\alpha}}$ (2.2)
and act on the vacuum states $|0\rangle$ and $\langle 0|$, $\langle 0|0\rangle=1$ in the standard way:
$a_{\alpha}|0\rangle=\bar{a}^{\dot{\alpha}}|0\rangle=0\,,\qquad\langle 0|\,c^{\alpha}=\langle 0|\,\bar{c}_{\dot{\alpha}}=0\,.$ (2.3)
It was shown in [12] that the Lagrangian dynamics of the infinite spin field in $\rm{AdS_{4}}$ space is described by the vector (2.1) satisfying the conditions
$L_{0}|{\Psi}\rangle=0\,,\qquad L|{\Psi}\rangle=0\,,\qquad L^{+}|{\Psi}\rangle=0\,,\qquad U|{\Psi}\rangle=0\,,$ (2.4)
with the constraint operators
@@
where $g=\det g_{\mu\nu}$. Relation (2.12) allows one to call the operator $(D)^{2}$ the generalized Laplace-Beltrami operator in the Fock space. Additionally, in (2.6) and (2.7), we use the non-indexed notation: $(c\sigma^{m}\bar{c}):=c^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}c^{\dot{\alpha}}$, $(a\sigma^{m}\bar{a}):=a^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}a^{\dot{\alpha}}$.
@@
First, the operator $U$ in the last condition (2.4) is the $\mathrm{U}(1)$ symmetry generator. The solution of the equation $U|{\Psi}\rangle=0$ is the vector (2.1), where $s_{1}$ and $s_{2}$ are equal to one another, $s_{1}=s_{2}=s$, and the expansion contains only the fields $\psi_{\alpha_{1}\ldots\alpha_{s}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s}}(x)$ with the equal number of undotted and dotted spinor indices. These are exactly the states, resolving the constraints $U|{\Psi}\rangle=0$, that were used in constructing the Lagrangian formulation in [12].
Second, the limit $R\to\infty$ (together with $\omega_{\mu}{}^{kl}=0$ and $e_{m}^{\mu}=\delta_{m}^{\mu}$) is well defined: it simply switches off the external gravitational field. In this limit, equations (2.4) yield the conditions that define the well-known representation of the Poincáre group with infinite spin in flat space. Indeed, in the flat limit the operators (2.5)-(2.7) take the form
@@
Taking into account relations (2.14), one gets that the condition $W^{m}W_{m}=\boldsymbol{\mu}^{2}$ holds for vectors obeying the first three equations (2.4) in the flat case $R\to\infty$. The last condition in (2.8) removes the degeneracy arising from the use of additional spinor variables in our approach.
Thus, in the flat case $R\to\infty$, the conditions (2.4) define the infinite spin field. On the other hand, equations (2.4) for a nonzero curvature $\kappa=-1/R^{2}\neq 0$ provide a consistent dynamical description of such a system in $AdS_{4}$ space.
@@ 3 Geometry of $\rm{AdS_{4}}$ space
in five-dimensional two-time space $\mathbb{R}^{2,3}$ with the metric $\eta_{MN}=\mathrm{diag}(-,+,+,+,-)$ and the coordinates $X^{M}$, $M=0,1,2,3,4$. The manifold defined by the condition (3.1) is the coset $\mathrm{SO}(2,3)/\mathrm{SO}(1,3)$. The symmetry generators of $\rm{AdS_{4}}$ space are 666 Throughout this paper we use the anti-Hermitean form of generators: for example, $(J_{MN})^{\dagger}={}-J_{MN}$.
@@
In the curve variable $x^{\mu}$ in $\rm{AdS_{4}}$ space (3.1) the generators $J_{mn}$ defined in (3.2) are represented only by orbital generators $J_{mn}=\mathcal{L}_{mn}$ which have the form (here $\eta_{m\mu}=\mathrm{diag}(-,+,+,+)$):
@@
The operators (3.15) and (3.13) realize the representation of the $\rm{AdS_{4}}$ algebra (3.5) in the space of scalar functions depending on the curve coordinate $x^{\mu}$. However, the infinite spin system under consideration is defined by conditions (2.4) on the Fock space vector (2.2), (2.3), where the creation $c^{\alpha}$, $\bar{c}_{\dot{\alpha}}$ and annihilation $a_{\alpha}$, $\bar{a}^{\dot{\alpha}}$ operators act, carrying indices of the tangent space structural group $\mathrm{SO}(1,3)$. This additional structure is taken into account by the replacement $\mathcal{L}_{mn}\ \to\ \mathcal{L}_{mn}+\mathcal{M}_{mn}$ in the operators (3.15) and (3.13) where $\mathcal{M}_{mn}$ are defined in (2.11).
@@ 4 $\mathrm{SO}(2,3)$ symmetry of infinite spin field
we can be represent the operator $L_{0}$ (2.5) in the form:
@@ 5.3 Irreducible representation
From the first condition $L_{0}|{\Psi}\rangle=0$ of the set (2.4) we obtain that the infinite spin vector (2.1) is the eigenvector of the second-order Casimir operator
$\mathcal{C}_{2}|{\Psi}\rangle=c_{2}|{\Psi}\rangle\,$ (5.37)
@@
Besides, due to the first three conditions in (2.4), namely, the conditions $L_{0}|{\Psi}\rangle=0$, $L|{\Psi}\rangle=0$, $L^{+}|{\Psi}\rangle=0$, the same vector $|{\Psi}\rangle$ is also the eigenvector of the fourth order Casimir operator
$\mathcal{C}_{4}|{\Psi}\rangle=c_{4}|{\Psi}\rangle$ (5.39)
@@
Thus, the vectors (2.1) under the conditions (2.4) are the eigenvectors of all Casimir operators of the $\mathfrak{so}(2,3)$ algebra with the continuous eigenvalues (5.38) and (5.40). Therefore, the vector $|{\Psi}\rangle$ describing the infinite spin field dynamics in $\rm{AdS_{4}}$ space corresponds to the continuous irreducible representation of the $\mathrm{SO}(2,3)$ group. The last condition in (2.4) removes the degeneracy of the spectrum associated with the use of additional spinor operators.

### v3

@@ 1 Introduction
One of the relatively recent trends in higher spin field theory [^1] Current state of the higher spin field problem is considered, e.g., in the reviews [1, 2, 3]. is the study of the Lagrangian formulation of infinite (continuous) field models in $\rm{AdS}$ space. Various approaches to this formulation are discussed in refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. [^2] The review of the infinite (continuous) spin field models is given in ref. [14], for more recent references see e.g., [12]. In particular, a full classification of possible unitary and non-unitary continuous spin field models was given in the paper by R. Metsaev [9] within the light-cone formalism.
@@
In a recent paper [12], a completely consistent covariant Lagrange formulation for bosonic infinite spin field in $\rm{AdS}_{4}$ space was constructed in the framework of the BRST approach to higher spin fields. The formulation is realized in Fock space vectors $|{\Psi}\rangle$ where the BRST charge is constructed based on a set of constraint operators satisfying the closed algebra (see the operators (2.5)-(2.8) bellow). The Lagrangian includes the real continuous parameter $\boldsymbol{\mu}$ and in the flat limit this Lagrangian is reduced to the known infinite spin field Lagrangian in Minkowski space [15] in which $\boldsymbol{\mu}$ defines the corresponding Poincaré representation. The on-shell structure of the theory under consideration can also be formulated in terms of the field $\Phi$ which depends on additional spinor variables and obeys all the same constraints as the formulation in the Fock space (see (2.4 below) encoded in the BRST charge. The relation between the descriptions in terms of Fock space vectors $|{\Psi}\rangle$ and on-shell fields $\Phi$ was discussed in detail in [15] for infinite (continuous) spin representations in flat space. Then natural questions arise as to what representation of the symmetry group of $\rm{AdS}$ space such a formulation corresponds and how it relates to Metsaev’s classification. In this paper, we answer these questions. First, we show that the field that satisfies the system of constraints obtained within the Lagrangian formulation in the paper [12] corresponds to the most degenerate continuous unitary representation of the Lie algebra $\mathfrak{so}(2,3)$. [^3] The notion of the most degenerate representation is based on the Gelfand theorem [16] which states that the number of independent invariant operators in the enveloping algebra acting in the Hilbert space of functions on the domain $X$ can be independent of the rank of the group. The most degenerate representation, if it exists in given space, means that the number of these operators is one. The construction of the most degenerate continuous representations of the group $\mathrm{SO}(p,q)$ is discussed in the papers [17, 18, 19], see also [20]. Second, we show that the representations of the $\mathrm{SO}(2,3)$ group realized on such a field correspond to the $\mathrm{SO}(2,3)$ representations in Metsaev’s classification. In this case, the eigenvalues of the $\mathfrak{so}(2,3)$ Casimir operators are determined by the dimensionless quantity $\boldsymbol{\mu}R$ where $R$ is the radius of $\rm{AdS}_{4}$ space.
@@
The paper is organized as follows. In Section 2, we briefly consider a reformulation in terms of the on-shell field $\Phi$ the results obtained in $AdS_{4}$ space [12], similar to what was done in flat space [15]. Here the on-shell infinite spin field in $\rm{AdS}_{4}$ space is defined, the operator constraints $L_{0},L,L^{+}$ are introduced, their algebra is written down and the flat limit is discussed. In Section 3, the geometric structure of $\rm{AdS}_{4}$ space is described in a form convenient for further consideration. In Section 4, we show that the operator constrains are invariants of the $\mathrm{SO}(2,3)$ group and one more invariant operator $(\Pi)^{2}$ is introduced. In Section 5, we construct the Casimir operators $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ of the $\mathfrak{so}(2,3)$ algebra acting in the space of fields $\Phi$ under the conditions (2.4). In doing so, we obtain two important statements. First, we prove that the Casimir operators $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ are completely expressed in terms of the operators of the constraints (2.4) under consideration. Second, we show that only one of these two Casimir operators is independent. This exactly corresponds to the definition of the most degenerate representation for the $\mathrm{SO}(2,3)$ group. The summary is a collection of results.
@@ 2 Infinite spin field in $\rm{AdS_{4}}$ space
We begin with a brief summary of the basic aspects related to the Lagrangian formulation of the infinite spin field in $\rm{AdS_{4}}$ space .
A consistent Lagrangian formulation for an infinite spin field in flat space was constructed in paper [15]. Following this paper, generalization for the infinite spin theory in $\rm{AdS_{4}}$ space is conveniently realized in terms of fields of the form
${\Phi}={\Phi}(x,c,\bar{c})\,,$ (2.1)
where $x^{\mu}$, $\mu=0,1,2,3$ are the coordinates on $\rm{AdS_{4}}$ space and $c^{\alpha}$, $\bar{c}_{\dot{\alpha}}=\epsilon_{\dot{\alpha}\dot{\beta}}\bar{c}^{\dot{\beta}}$, $\bar{c}^{\dot{\alpha}}=(c^{\alpha})^{\dagger}$ are the components of the commuting Weyl spinor. [^4] The two-component Weyl spinor indices are raised and lowered by $\epsilon_{\alpha\beta}=-\epsilon^{\alpha\beta}$, $\epsilon_{\dot{\alpha}\dot{\beta}}=-\epsilon^{\dot{\alpha}\dot{\beta}}$: $\varphi^{\alpha}=\epsilon^{\alpha\beta}\varphi_{\beta}$, $\bar{\varphi}^{\dot{\alpha}}=\epsilon^{\dot{\alpha}\dot{\beta}}\bar{\varphi}_{\dot{\beta}}$. The relativistic Pauli matrices $\sigma^{m}_{\alpha\dot{\beta}}$ and $\tilde{\sigma}_{m}^{\dot{\alpha}\beta}$ satisfy the basic relation $\sigma_{m}\tilde{\sigma}_{n}+\sigma_{n}\tilde{\sigma}_{m}={}-2\eta_{mn}$ where $\eta_{mn}=\mathrm{diag}(-1,+1,+1,+1)$. We use the matrices $\sigma_{mn}=-\left(\sigma_{m}\tilde{\sigma}_{n}-\sigma_{n}\tilde{\sigma}_{m}\right)/4$, $\tilde{\sigma}_{mn}=-\left(\tilde{\sigma}_{m}\sigma_{n}-\tilde{\sigma}_{n}\sigma_{m}\right)/4$. Here and further the Latin indices take the values 0,1,2,3. The construction of fields (2.1) is discussed in details in [15]. Also we use the operators
$a_{\alpha}=\frac{\partial}{\partial c^{\alpha}}\,,\qquad\bar{a}^{\dot{\alpha}}=\frac{\partial}{\partial\bar{c}_{\dot{\alpha}}}\,,$ (2.2)
satisfying the following commutation relations:
$[a_{\alpha},c^{\beta}]=\delta_{\alpha}^{\beta}\,,\qquad[\bar{a}^{\dot{\alpha}},\bar{c}_{\dot{\beta}}]=\delta_{\dot{\beta}}^{\dot{\alpha}}\,.$ (2.3)
The infinite spin states in $\rm{AdS_{4}}$ are described by the fields (2.1) subject to the conditions
$L_{0}{\Phi}=0\,,\qquad L{\Phi}=0\,,\qquad L^{+}{\Phi}=0\,,\qquad U{\Phi}=0\,,$ (2.4)
with the constraint operators found in [12] in the form
@@
where $g=\det g_{\mu\nu}$. Relation (2.12) allows one to call the operator $(D)^{2}$ the generalized Laplace-Beltrami operator in space of vectors (2.1). Additionally, in (2.6) and (2.7), we use the non-indexed notation: $(c\sigma^{m}\bar{c}):=c^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}c^{\dot{\alpha}}$, $(a\sigma^{m}\bar{a}):=a^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}a^{\dot{\alpha}}$.
@@
First, the operator $U$ in the last condition (2.4) is the $\mathrm{U}(1)$ symmetry generator [12]. Second, the limit $R\to\infty$ (together with $\omega_{\mu}{}^{kl}=0$ and $e_{m}^{\mu}=\delta_{m}^{\mu}$) is well defined: it simply switches off the external gravitational field. In this limit, equations (2.4) yield the conditions that define the well-known representation of the Poincáre group with infinite spin in flat space. Indeed, in the flat limit the operators (2.5)-(2.7) take the form
@@
Taking into account relations (2.14), one gets that the condition $W^{m}W_{m}=\boldsymbol{\mu}^{2}$ holds for functions obeying the first three equations (2.4) in the flat case $R\to\infty$. The last condition in (2.8) removes the degeneracy arising from the use of additional spinor variables in our approach.
Thus, in the flat case $R\to\infty$, the conditions (2.4) define the infinite spin field.  [^6] Note that our formulation of infinite spin fields uses spinor coordinates as additional variables unlike the vector additional coordunates in the seminal papers [21, 22] (see also [14] for review). Therefore, although our description of the field differs from the Bargman-Wigner formulation, the correct value of the squared Pauli-Lubanski vector (2.15) guarantees a correct description of massless irreducible representations of infinite (continuous) spin in flat space (see discussion after (2.15)). Besides, it was shown in [23] that the Bargman-Wigner formulation with vector additional coordinates is reproduced within the formulation with spinor additional coordinates if to carry out a spinor integral transform. We believe that the description using spinor additional coordinates is technically simpler. On the other hand, equations (2.4) for a nonzero curvature $\kappa=-1/R^{2}\neq 0$ provide a consistent dynamical description of such a system in $AdS_{4}$ space.
@@ 3 Geometry of $\rm{AdS_{4}}$ space
in five-dimensional two-time space $\mathbb{R}^{2,3}$ with the metric $\eta_{MN}=\mathrm{diag}(-,+,+,+,-)$ and the coordinates $X^{M}$, $M=0,1,2,3,4$. The manifold defined by the condition (3.1) is the coset $\mathrm{SO}(2,3)/\mathrm{SO}(1,3)$. The symmetry generators of $\rm{AdS_{4}}$ space are 777 Throughout this paper we use the anti-Hermitean form of generators: for example, $(J_{MN})^{\dagger}={}-J_{MN}$.
@@
In the curved variable $x^{\mu}$ in $\rm{AdS_{4}}$ space (3.1) the generators $J_{mn}$ defined in (3.2) are represented only by orbital generators $J_{mn}=\mathcal{L}_{mn}$ which have the form (here $\eta_{m\mu}=\mathrm{diag}(-,+,+,+)$):
@@
The operators (3.15) and (3.13) realize the representation of the $\rm{AdS_{4}}$ algebra (3.5) in the space of scalar functions depending on the curved coordinate $x^{\mu}$. However, the infinite spin system under consideration is defined by conditions (2.4) on the Fock space vector (2.3), (2.2), where the creation $c^{\alpha}$, $\bar{c}_{\dot{\alpha}}$ and annihilation $a_{\alpha}$, $\bar{a}^{\dot{\alpha}}$ operators act, carrying indices of the tangent space structural group $\mathrm{SO}(1,3)$. This additional structure is taken into account by the replacement $\mathcal{L}_{mn}\ \to\ \mathcal{L}_{mn}+\mathcal{M}_{mn}$ in the operators (3.15) and (3.13) where $\mathcal{M}_{mn}$ are defined in (2.11).
@@ 4 $\mathrm{SO}(2,3)$ symmetry of infinite spin field
we can represent the operator $L_{0}$ (2.5) in the form:
@@ 5.3 Irreducible representation
From the first condition $L_{0}{\Phi}=0$ of the set (2.4) we obtain that the infinite spin field (2.1) is the eigenvector of the second-order Casimir operator
$\mathcal{C}_{2}{\Phi}=c_{2}{\Phi}\,$ (5.37)
@@
Besides, due to the first three conditions in (2.4), namely, the conditions $L_{0}{\Phi}=0$, $L{\Phi}=0$, $L^{+}{\Phi}=0$, the same field ${\Phi}$ is also the eigenvector of the fourth order Casimir operator
$\mathcal{C}_{4}{\Phi}=c_{4}{\Phi}$ (5.39)
@@
Thus, the field (2.1) under the conditions (2.4) is the eigenvector of all Casimir operators of the $\mathfrak{so}(2,3)$ algebra with the continuous eigenvalues (5.38) and (5.40). Therefore, the field $\Phi$ describing the infinite spin field dynamics in $\rm{AdS_{4}}$ space corresponds to the continuous irreducible representation of the $\mathrm{SO}(2,3)$ group. The last condition in (2.4) removes the degeneracy of the spectrum associated with the use of additional spinor operators.